OpenStudy (anonymous):

Please help! (6x-5)=(x+10) How do I work this out step by step?

OpenStudy (anonymous):

First thing we can do is get rid of parenthesis. \[ 6x-5=x+10 \]

OpenStudy (anonymous):

Alright, that's actually the first step I have down!

OpenStudy (anonymous):

Next thing we want to is subtract \(x\) from both sides.\[ (6x-5)-x=(x+10)-x \]

OpenStudy (anonymous):

Again, the parenthesis can be removed.\[ 6x-5-x=x+10-x \]Do you know how to combine like terms?

OpenStudy (anonymous):

The \(6x\) and \(-x\) combine into \(5x\), and the \(x\) and \(-x\) cancel out into \(0\).

OpenStudy (anonymous):

Which gives us: \[ 5x-5=10 \]

OpenStudy (anonymous):

I think I can work it out from there, thank you so much!

OpenStudy (anonymous):

You are welcome, I guess.